# Benchmarking stabilized and self-stabilized $p$-virtual element methods with variable coefficients

### v1

@@ Abstract
Standard Virtual Element Methods (VEM) are based on polynomial projections and require a stabilization term to evaluate the contribution of the non-polynomial component of the discrete space. However, the stabilization term does not generally conform to the physics of the considered problem and a criterion for its choice is not typically supported by theoretical arguments. Hence, stabilization-free and self-stabilized formulations have been proposed. Moreover, the accuracy of VEM deteriorates in case of problems with variable coefficients, that are not usually accounted when constructing polynomial projections. The mentioned aspects might limit the use of the method when tackling real-world applications.
This paper provides an in-depth numerical investigation into different stabilized and self-stabilized formulations for the $p$-version of VEM. The results show that self-stabilized and stabilization-free formulations achieve optimal accuracy, while suffering from worse conditioning. Moreover, a new approach for dealing with variable coefficients is introduced: the discrete space is modified so that general coefficients can be accounted when computing the polynomial projectors. Numerical results show that this new approach is more robust than the standard ones.
@@ 1 Introduction
Current research studies focus on the design of novel aerospace structures with improved efficiency. An example is given by variable stiffness panels, which are obtained by combining a variable stiffness skin, in which the fibers’ orientation varies across the domain, with stringers running along curvilinear paths. These configurations are modeled by partial differential equations with non-uniform elastic properties over the panel domain. Numerical approaches for the mentioned application have been proposed in the framework of finite elements [zhao2019prestressed] and Ritz method [vescovini2020semi]. However, in a finite element context, node sharing between plate and stiffeners requires careful meshing or ad-hoc formulations [zhao2017vibration]. On the contrary, a Ritz approach simplifies the enforcement of the compatibility between plate and stringers, but is generally not versatile in dealing with complex geometries, unless specific procedures are employed [vescovini2023ritz].
@@
In this landscape, the Virtual Element Method (VEM) [daveiga2013basic] is a promising tool since it supports the use of meshes made of general polytopal elements, with possibly curved edges [daveiga2019virtual2]. Thus, meshes conforming with the stringers can be easily generated and combined with high-order discretizations.
@@
Despite VEM has already been applied to several problems in computational mechanics, such as plane stress elasticity [daveiga2013virtual, artioli2017arbitrary, artioli2017stress, mengolini2019engineering, daltri2020error], buckling problems [mora2020virtual, meng2020linear, adak2023c0], Kirchhoff–Love [brezzi2013virtual] and Reissner–Mindlin plate models [daveiga2019virtual, daltri2022first], to the best of the authors’ knowledge its application to variable stiffness panels remains limited. As a first step in this direction, the present paper serves as a preliminary investigation to benchmark virtual elements in their $p$-version [daveiga2016basic] by taking into account the challenges posed by variable stiffness panels.
@@
VEM can be interpreted as a generalization of finite elements on polytopal meshes. However, virtual elements give rise to a nonconforming method as the trial and test functions are solutions of local PDEs (defined in each mesh element) that are never explicitly solved when constructing the discrete formulation. Hence the name “Virtual”. Since only a polynomial projection of the trial functions is computable through the degrees of freedom, a stabilization term is introduced to handle the residual non-polynomial component and ensure the stability of the discrete problem. However, the stabilization term has an arbitrary nature and does not generally conform to the physics of the problem. In the particular case of variable coefficients, this term is scaled by a constant approximation and the polynomial projection is built either independently of the coefficient [daveiga2016virtual] or by taking into account a low-order polynomial approximation thereof [daveiga2014virtual]. Moreover, the stabilization involves the introduction of a user-defined scaling parameter, which reduces the generality of the method, potentially leading to over-stabilization and degraded solution quality: this situation is verified for anisotropic problems [berrone2022comparison] and eigenvalue problems [boffi2020approximation, alzaben2025stabilization], where an improper choice of the stability parameters may trigger the onset of spurious modes. The relation between the stabilization parameter and the performance of VEM has been discussed in [cangiani2015hourglass, russo2022quantitative]. Despite the many stabilization forms available in the literature [daveiga2013basic, daveiga2017stability, daveiga2018exponential, mascotto2018ill, bertoluzza2022stabilization], most of them are suitable for low-order virtual elements, while they tend to be suboptimal as the order increases.
@@
An approach to overcome the issues caused by the stabilization term is given by stabilization-free virtual elements [berrone2025lowest, berrone2022comparison, berrone2025stabilization]. These elements are obtained by projecting the virtual functions onto larger polynomial spaces. So, the resulting discrete problem is automatically stable. A different strategy refers to the so-called self-stabilized elements, which are obtained by enriching the approximation space with extra internal degrees of freedom [lamperti2023hu]. In this work, the term self-stabilized VEM will refer to both methods throughout this paper. Although not addressed here, the work in [credali2024reduced] presents a reduced basis approach to cheaply reconstruct the lowest-order VEM basis functions and avoid the need of the stabilization term.
@@
The second element of novelty of this paper regards a new approach for dealing with problems governed by variable coefficients in high-order approximations. The virtual element discrete space is enhanced by including the coefficients effect. This novel framework, denoted as $\mathrm{VC}$-VEM, allows the polynomial projections to be explicitly computed by incorporating the coefficient in their definition. This formulation can be easily embedded into stabilized and self-stabilized settings, leading to improved accuracy.
@@ 2 Geometrical framework
The domain is denoted by $\Omega$ and $\left\{\Gamma_{i}\right\}_{i=1,\dots,N_{e}}$ is a finite set of smooth curves identifying the boundary $\Gamma:=\partial\Omega$, where $N_{e}$ is the number of edges. By assuming that the boundary $\Gamma$ is Lipschitz and that for $m\geq 0$ each curve $\Gamma_{i}$ is of class $\mathcal{C}^{m+1}$ [daveiga2019virtual2], one can define the regular and invertible $\mathcal{C}^{m+1}$ parametrization of $\Gamma_{i}$ with respect to $I_{i}=\left[a_{i},b_{i}\right]$ as $\gamma_{i}:I_{i}\rightarrow\Gamma_{i}$. As $\gamma_{i}$ is an invertible mapping, for a generic function $g$:
@@
The domain $\Omega$ is partitioned by a finite set $\mathcal{T}_{h}$ of non-overlapping star-shaped polygons with possibly curved edges $E\in\mathcal{T}_{h}$, as in the standard VEM theory [daveiga2013basic]. The symbol $h_{E}$ denotes the diameter of $E$ and $h=\max_{E\in\mathcal{T}_{h}}h_{E}$. The boundary of $E$ is denoted by $\Gamma^{E}=\partial E$. More precisely, $\Gamma^{E}$ is the union of straight segments $\Gamma^{E}_{e}$ with $e=1,\dots,n_{e}$ and curved edges $\tilde{\Gamma}^{E}_{e}$ with $e=1,\dots,\tilde{n}_{e}$, satisfying the regularity assumptions mentioned above.
@@ 3.1 Stabilized VEM
The accuracy of the method is guaranteed by the inclusion $\mathcal{P}_{p}\left(E\right)\subset\mathcal{V}_{h}\left(E\right)$ and the operator $\Pi^{\nabla}_{p}$ maps the discrete functions from the VEM space to the polynomial one and will be defined later. Conventionally, $\mathcal{P}_{-1}\left(E\right)=\left\{0\right\}$.
Moreover, the polynomial space $\tilde{\mathcal{P}}_{p}\left(\tilde{\Gamma}^{E}_{e}\right)$ on the curved edges is defined by using Eq. (1) [daveiga2019virtual2]:
@@
    $\textbf{DOF}_{2}$: for $p>1$, the values at the p1p-1 internal Gauss-Lobatto quadrature points on each $\Gamma^{E}_{e}$ and the values of at the p1p-1 points on each $\tilde{\Gamma}^{E}_{e}$ that are images through $\gamma_{e}$ of the p1p-1 internal Gauss-Lobatto quadrature points on $I_{e}$ [daveiga2019virtual2],
@@
$+\sum_{e=1}^{\tilde{n}_{e}}\int_{\tilde{\Gamma}^{E}_{e}}\psi_{i}\cdot\nabla q\,\bm{\hat{n}}_{\tilde{\Gamma}_{e}}\;\,\mathrm{d}\tilde{\Gamma}^{E}_{e}$
The surface integral of the equation above is computed from the internal degrees of freedom, while the line integrals on the straight edges are evaluated by standard quadrature rules for polynomials. On the other hand, the discrete functions $\psi_{i}$ are not polynomials on curved edges, so that the parametrization $\gamma_{e}$ defined in Eq. (1) is used to recover the polynomial structure on the reference interval $I_{e}$ [daveiga2019virtual2]:
@@
$=\sum_{e=1}^{\tilde{n}_{e}}\int_{I_{e}}\left[\tilde{\psi_{i}}\cdot\nabla\tilde{q}\;\tilde{\bm{\hat{n}}}_{\tilde{\Gamma}_{e}}\right]\left\lVert\gamma_{e}^{{}^{\prime}}\right\rVert\,\mathrm{d}I_{e}$
where $\tilde{\psi_{i}}$ is the image of $\psi_{i}$ through $\gamma_{e}$. Once the projection $\Pi^{\nabla}_{p}$ is computed, the VEM space is decomposed into a polynomial space and its orthogonal counterpart so that:
@@
By leveraging the above orthogonal decomposition, the bilinear form reads:
@@
$\alpha_{*}a^{E}\left(\delta u_{h},\delta u_{h}\right)\leq S^{E}\left(\delta u_{h},\delta u_{h}\right)\leq\alpha^{*}a^{E}\left(\delta u_{h},\delta u_{h}\right),$ (19)
@@
where $\omega_{r}=\max\left\{1,a^{E}\left(\psi_{r},\psi_{r}\right)\right\}$. Observe that $S^{E}_{2}$ only accounts for the boundary degrees of freedom, whereas $S^{E}_{3}$ scales as the consistency part, making it more robust than $S^{E}_{1}$ and $S^{E}_{2}$.
@@
The last stabilization considered here refers to [artioli2017arbitrary, daltri2022first] and is particularly effective in linear elasticity problems. It is usually expressed in matrix form as:
@@ 3.2 Self-stabilized VEM
As detailed in the previous section, the arbitrary nature of stabilization terms motivates the seek for more general formulations, not relying upon heuristic terms. Thus, self-stabilized formulations have been proposed as a viable strategy to prevent the used of the stabilization term. These approaches are particularly useful when dealing with high-order approximations and more complex problems.
@@
V1 Berrone et al. [berrone2025lowest] Enlarged enhanced
V2 Lamperti et al. [lamperti2023hu] Augmented Additional dofs
V3 Berrone et al. [berrone2025stabilization] Standard Div-free polynomials
V4 inspired by [berrone2025lowest] Enlarged enhanced
V5 inspired by [lamperti2023hu] Augmented Additional dofs
V6 inspired by [berrone2025stabilization] Standard Div-free polynomials
@@
The first self-stabilized technique was introduced in [berrone2025lowest] for the lowest-order VEM and here extended to the generic order $p$.
@@
This formulation refers to the one presented in [lamperti2023hu] for the lowest-order VEM for linear elasticity problems and here extended to generic order.
The augmented space is defined as:
@@
Note, if $\ell_{E}=1$, from Eq. (29) one obtains that $\Delta u_{h}\big|_{E}\in\mathcal{P}_{p-1}\left(E\right)$. Hence, the definition of the space is modified to guarantee the computability of the required quantities:
@@
This results in adding internal moments of order p1p-1 in combination with the standard enhanced VEM condition for computing the $L^{2}$ product with polynomials of degree $p$.
@@
The third self-stabilized formulation is taken from [berrone2025stabilization] and refers to the same VEM space of the classical stabilized VEM. The polynomial projection is then computed onto the space
@@
for all $i=1,\dots,N_{d}\big|_{E}$. The divergence-free property is convenient since the surface integral vanishes after integrating by parts $\left(\nabla\psi_{i},\text{curl}\left(q\right)\right)_{E}$.
@@
In this case, the VEM space is the standard one, defined in Eq. (8), while the polynomial projection is computed in the space
@@ 3.2.1 Algorithm for computing $\ell_{E}$ on a given element $E$
The choice of $\ell_{E}$ is crucial for the definition of the proposed schemes. The following procedure applies to all the aforementioned self-stabilized formulations. The following necessary condition must be satisfied [berrone2025stabilization]:
@@
The inequality of Eq. (38) is necessary but not sufficient. The order $p+\ell_{E}$ will be chosen so that the local stiffness matrix $\bm{K}^{E}$ is full rank. The order $\ell_{E}$ can be determined by applying the procedure outlined in Algorithm 1, which differs from the one introduced in [berrone2025lowest], in which a QR decomposition is employed. Notice that Algorithm 1 is preferred solely for simplicity of implementation.

### v2

@@ Abstract
Standard Virtual Element Methods (VEM) are based on polynomial projections and require a stabilization term to evaluate the contribution of the non-polynomial component of the discrete space. However, the stabilization term is not uniquely defined by the underlying variational formulation and is typically introduced in an ad hoc manner, potentially affecting the numerical response. Stabilization-free and self-stabilized formulations have been proposed to overcome this issue, although their theoretical analysis is still less mature.
This paper provides an in-depth numerical investigation into different stabilized and self-stabilized formulations for the $p$-version of VEM. The results show that self-stabilized and stabilization-free formulations achieve optimal accuracy while suffering from worse conditioning. Moreover, a new projection operator, which explicitly accounts for variable coefficients, is introduced within the framework of standard virtual element spaces. Numerical results show that this new approach is more robust than the existing ones for large values of $p$.
@@ 1 Introduction
Current research studies focus on the design of novel aerospace structures with improved efficiency. An example is given by variable stiffness panels, which are obtained by combining a variable stiffness skin, in which the fibers’ orientation varies across the domain, with stringers running along curvilinear paths. These configurations are modeled by partial differential equations with non-uniform elastic properties over the panel domain. Numerical approaches for the mentioned application have been proposed in the framework of finite elements [37] and Ritz method [34]. However, in a finite element context, node sharing between plate and stiffeners requires careful meshing or ad-hoc formulations [38]. On the contrary, a Ritz approach simplifies the enforcement of the compatibility between plate and stringers, but is generally not versatile in dealing with complex geometries, unless specific procedures are employed [35].
@@
In this landscape, the Virtual Element Method (VEM) [6] is a promising tool since it supports the use of meshes made of general polytopal elements, with possibly curved edges [14]. Thus, meshes conforming with the stringers can be easily generated and combined with high-order discretizations.
@@
Despite VEM has already been applied to several problems in computational mechanics, such as plane stress elasticity [8, 3, 4, 30, 24], buckling problems [31, 29, 1], Kirchhoff–Love [21] and Reissner–Mindlin plate models [13, 25], to the best of the authors’ knowledge its application to variable stiffness panels remains limited. A contribution can be found in [32], in which a piecewise constant approximation of the variable coefficient is considered. For elasticity problems, spatially varying material coefficients have been the subject of other investigations in a VEM framework based on the Hellinger–Reissner variational principle. In the two-dimensional setting of [5], a local energy projection operator is defined, leading to elementwise approximation spaces that are no longer polynomial in the presence of variable elasticity coefficients. A similar approach is proposed in [36]. The formulation is here three-dimensional and an analogous projection-based approach is used for polyhedral meshes. In this context, the present paper serves as a preliminary investigation to benchmark virtual elements in their $p$-version [9] by taking into account the challenges posed by variable stiffness panels.
@@
VEM can be interpreted as a generalization of finite elements on polytopal meshes. However, virtual elements give rise to a nonconforming method as the trial and test functions are solutions of local PDEs (defined in each mesh element) that are never explicitly solved when constructing the discrete formulation. Hence the name “Virtual”. Since only a polynomial projection of the trial functions is computable through the degrees of freedom, a stabilization term is introduced to handle the residual non-polynomial component and ensure the stability of the discrete problem. However, the stabilization term has an arbitrary nature and does not generally conform to the physics of the problem. In the particular case of variable coefficients, this term is scaled by a constant approximation and the polynomial projection is built either independently of the coefficient [7] or by taking into account a low-order polynomial approximation thereof [15].
The stabilization involves the introduction of a user-defined scaling parameter, which reduces the generality of the method, potentially leading to over-stabilization and degraded solution quality: this situation is verified for anisotropic problems [17] and eigenvalue problems [20, 2], where an improper choice of the stability parameters may trigger the onset of spurious modes. The relation between the stabilization parameter and the performance of VEM has been discussed in [22, 33]. Despite the many stabilization forms available in the literature [6, 11, 10, 28, 19], most of them are suitable for low-order virtual elements, while they tend to be suboptimal as the order increases.
@@
An approach to overcome the issues caused by the stabilization term is given by stabilization-free virtual elements [18, 17, 16]. These elements are obtained by projecting the virtual functions onto larger polynomial spaces. So, the resulting discrete problem is automatically stable. A different strategy refers to the so-called self-stabilized elements, which are obtained by enriching the approximation space with extra internal degrees of freedom [27]. The term self-stabilized VEM will refer to both methods throughout this paper. Although not addressed here, the work in [23] presents a reduced basis approach to cheaply reconstruct the lowest-order VEM basis functions and avoid the need of the stabilization term.
@@
The second element of novelty of this paper regards a new approach for dealing with problems governed by variable coefficients in high-order approximations. Within the framework of well-known virtual element spaces, a new polynomial projection is introduced so that the features of the coefficient are taken into account. This formulation, denoted as $\mathrm{VC}$-VEM, can be easily embedded into stabilized and self-stabilized settings, leading to improved accuracy.
@@ 2 Geometrical framework
The domain is denoted by $\Omega$ and $\left\{\Gamma_{i}\right\}_{i=1,\dots,N_{e}}$ is a finite set of smooth curves identifying the boundary $\Gamma:=\partial\Omega$, where $N_{e}$ is the number of edges. By assuming that the boundary $\Gamma$ is Lipschitz and that for $m\geq 0$ each curve $\Gamma_{i}$ is of class $\mathcal{C}^{m+1}$ [14], one can define the regular and invertible $\mathcal{C}^{m+1}$ parametrization of $\Gamma_{i}$ with respect to $I_{i}=\left[a_{i},b_{i}\right]$ as $\gamma_{i}:I_{i}\rightarrow\Gamma_{i}$. As $\gamma_{i}$ is an invertible mapping, for a generic function $g$:
@@
The domain $\Omega$ is partitioned by a finite set $\mathcal{T}_{h}$ of non-overlapping star-shaped polygons with possibly curved edges $E\in\mathcal{T}_{h}$, as in the standard VEM theory [6]. The symbol $h_{E}$ denotes the diameter of $E$ and $h=\max_{E\in\mathcal{T}_{h}}h_{E}$. The boundary of $E$ is denoted by $\Gamma^{E}=\partial E$. More precisely, $\Gamma^{E}$ is the union of straight segments $\Gamma^{E}_{e}$ with $e=1,\dots,n_{e}$ and curved edges $\tilde{\Gamma}^{E}_{e}$ with $e=1,\dots,\tilde{n}_{e}$, satisfying the regularity assumptions mentioned above.
@@ 3.1 Stabilized VEM
When $\tilde{n}_{e}=0$, i.e. $E$ has only straight edges, the inclusion $\mathcal{P}_{p}\left(E\right)\subset\mathcal{V}_{h}\left(E\right)$} guarantees the accuracy of the method. On the other hand, for $\tilde{n}_{e}>0$, $\mathcal{P}_{p}\left(E\right)\not\subset\mathcal{V}_{h}\left(E\right)$} and only $\mathcal{P}_{0}\left(E\right)\subset\mathcal{V}_{h}\left(E\right)$} holds: in this case, the accuracy is guaranteed by the fact that curved edges are a feature of the domain and do not depend on $h$, see [14] for more details. The operator $\Pi^{\nabla}_{p}$ maps the discrete functions from the VEM space to the polynomial one and will be defined later. Conventionally, $\mathcal{P}_{-1}\left(E\right)=\left\{0\right\}$.
Moreover, the polynomial space $\tilde{\mathcal{P}}_{p}\left(\tilde{\Gamma}^{E}_{e}\right)$ on the curved edges is defined by using Eq. (1) [14]:
@@
    $\textbf{DOF}_{2}$: for $p>1$, the values at the p1p-1 internal Gauss-Lobatto quadrature points on each $\Gamma^{E}_{e}$ and the values of at the p1p-1 points on each $\tilde{\Gamma}^{E}_{e}$ that are images through $\gamma_{e}$ of the p1p-1 internal Gauss-Lobatto quadrature points on $I_{e}$ [14],
@@
$+\sum_{e=1}^{\tilde{n}_{e}}\int_{\tilde{\Gamma}^{E}_{e}}\psi_{i}\cdot\nabla q\,\bm{\hat{n}}_{\tilde{\Gamma}_{e}}\;\,\mathrm{d}\tilde{\Gamma}^{E}_{e},$
The volume integral of the equation above is computed from the internal degrees of freedom, while the boundary integrals on the straight edges are evaluated by standard quadrature rules for polynomials. On the other hand, the discrete functions $\psi_{i}$ are not polynomials on curved edges, so that the parametrization $\gamma_{e}$ defined in Eq. (1) is used to recover the polynomial structure on the reference interval $I_{e}$ [14]:
@@
$=\sum_{e=1}^{\tilde{n}_{e}}\int_{I_{e}}\left[\tilde{\psi_{i}}\cdot\nabla\tilde{q}\;\tilde{\bm{\hat{n}}}_{\tilde{\Gamma}_{e}}\right]\left\lVert\gamma_{e}^{{}^{\prime}}\right\rVert\,\mathrm{d}I_{e},$
where $\tilde{\psi_{i}}$ is the image of $\psi_{i}$ through $\gamma_{e}$. Once the projection $\Pi^{\nabla}_{p}$ is computed, the following decomposition holds:
@@
By leveraging the definition of $\Pi^{\nabla}_{p}$, the bilinear form reads:
@@
$\alpha_{*}(p)\,a^{E}\left(\delta u_{h},\delta u_{h}\right)\leq S^{E}\left(\delta u_{h},\delta u_{h}\right)\leq\alpha^{*}(p)\,a^{E}\left(\delta u_{h},\delta u_{h}\right),$ (19)
@@
where $\omega_{r}=\max\left\{1,a^{E}\left(\Pi^{\nabla}_{p}\psi_{r},\Pi^{\nabla}_{p}\psi_{r}\right)\right\}$. Observe that $S^{E}_{2}$ only accounts for the boundary degrees of freedom, whereas $S^{E}_{3}$ scales as the consistency part, making it more robust than $S^{E}_{1}$ and $S^{E}_{2}$.
@@
The last stabilization considered here refers to [3, 25] and is particularly effective in linear elasticity problems. It is usually expressed in matrix form as:
@@ 3.2 Self-stabilized VEM
As detailed in the previous section, the arbitrary nature of stabilization terms motivates the seek for more general formulations, not relying upon heuristic terms. Thus, self-stabilized formulations have been proposed as a viable strategy to prevent the use of the stabilization term. These approaches are particularly useful when dealing with high-order approximations and more complex problems.
@@
V1 Berrone et al. [18] Enlarged enhanced
V2 Lamperti et al. [27] Augmented Additional dofs
V3 Berrone et al. [16] Standard Div-free polynomials
V4 inspired by [18] Enlarged enhanced
V5 inspired by [27] Augmented Additional dofs
V6 inspired by [16] Standard Div-free polynomials
@@
The first self-stabilized technique was introduced in [18] for the lowest-order VEM and here extended to the generic order $p$.
@@
This formulation refers to the one presented in [27] for the lowest-order VEM for linear elasticity problems and here extended to generic order.
If the additional polynomial order is $\ell_{E}>1$, the augmented space is defined as:
@@
Note, if $\ell_{E}=1$, from Eq. (29) one obtains that $\Delta u_{h}\big|_{E}\in\mathcal{P}_{p-1}\left(E\right)$, which results in a weaker condition than that of the standard enhanced space in Eq. (8). Hence, for the particular case of $\ell_{E}=1$, the following variant of $\mathcal{V}^{V2}_{h}\left(E\right)$ is considered to guarantee the computability of the $L^{2}$ product with polynomials of degree $p$:
@@
As a consequence, internal moments of order p1p-1 are added in combination with the standard enhanced VEM condition for computing the $L^{2}$ product with polynomials of degree $p$.
@@
The third self-stabilized formulation is taken from [16] and refers to the same VEM space of the classical stabilized VEM. The polynomial projection is then computed onto the space:
@@
for all $i=1,\dots,N_{d}\big|_{E}$. The divergence-free property is convenient since the volume integral vanishes after integrating by parts $\left(\nabla\psi_{i},\text{curl}\left(q\right)\right)_{E}$.
@@
In this case, the VEM space is the standard one, defined in Eq. (8), while the polynomial projection is computed in the space:
@@ 3.2.1 Algorithm for computing $\ell_{E}$ on a given element $E$
The choice of $\ell_{E}$ is crucial for the definition of the proposed schemes. However, a general theory for the efficient choice of $\ell_{E}$ is still lacking; thus, an iterative procedure is carried out to determine its value [16]. The following procedure applies to all the aforementioned self-stabilized formulations. The following necessary condition must be satisfied [16]:
@@
The inequality of Eq. (38) is necessary but not sufficient. The order $p+\ell_{E}$ will be chosen so that the local stiffness matrix $\bm{K}^{E}$ is full rank. The order $\ell_{E}$ can be determined by applying the procedure outlined in Algorithm 1, which differs from the one introduced in [18], in which a QR decomposition is employed. Notice that Algorithm 1 is preferred solely for simplicity of implementation.